COPRODUCTS OF BOUNDED DISTRIBUTIVE LATTICES: CANCELLATION (A BANFF CONFERENCE PROBLEM)

JONATHAN DAVID FARLEY

Abstract. Let L ∗ M denote the coproduct of the bounded distributive lattices L and M. At the 1981 Banff Conference on Ordered Sets, the following question was posed: What is the largest class L of finite distributive lattices such that, for every non-trivial Boolean lattice B and every L ∈ L, B ∗ L = B ∗ L′ implies L = L′?

In this note, the problem is solved.